Show that 1n^3+2n+3n^2 is divided by 2 and 3 for all positive integers n
1
Expert's answer
2019-09-20T09:50:35-0400

Answer.  Define "f(n) = 1n^3+2n+3n^2". We try to factorize this polynomial. One obvious factor is "n", so "f(n) = n(1n^2+2+3n)". We factorize the factor of degree "2" by completing the square:


"1n^2+2+3n = \left(n+\frac{3}{2}\right)^2 - \frac{9}{4} +2 = \left(n+\frac{3}{2}\right)^2 - \left(\frac{1}{2}\right)^2"

"= \left(n +\frac{3}{2} -\frac{1}{2}\right) \left(n +\frac{3}{2} +\frac{1}{2}\right) = (n+1)(n+2)."

Hence

"f(n) = n (n+1) (n+2)."

Let "n" be a positive integer number. Division of "n" by "3" gives an integer quotient "q" and a remainder "r\in \{0, 1, 2\}" such that "n = 3q+r". Now check all the possible values of "r".

  • If "r=0", then "3" divides "n = 3q" which is a factor of "f(n)".
  • If "r=1", then "3" divides "n+2 = (3q+1) +2 = 3(q+1)" which is a factor of "f(n)".
  • If "r=2", then "3" divides "n+1 = (3q+2) +1 = 3(q+1)" which is a factor of "f(n)".

Hence "3" divides "f(n)" in all cases.

Division of "n" by "2" gives an integer quotient "q" and a remainder "r\in \{0, 1\}" such that "n = 2q+r".

  • If "r=0", then "2" divides "n = 2q" which is a factor of "f(n)".
  • If "r=1", then "2" divides "n+1 = (2q+1) +1 = 2(q+1)" which is a factor of "f(n)".

Hence "2" divides "f(n)" in all cases.